# Eulerian magnitude homology: diagonality, injective words, and regular path homology

### v1

@@ 5.2. The regular magnitude-path spectral sequence
Let $T$ be a finite directed tree, a bipartite graph with alternating orientation or a transitive tournament. Then, its (regular) path homology is trivial.
@@
We have seen that finite directed tree, bipartite graphs with alternating orientation or transitive tournaments are regularly diagonal. Furthermore, the associated directed flag complex is contractible. Hence, their (regular) path homology is trivial. ∎
As another example, consider the directed graphs of Example 4.8. They are both regularly diagonal and connected, on the same number of vertices. Their regular path homology is the homology of the $2$-sphere, as $\mathrm{dFl}\circ\mathrm{Pre}$ yields the $2$-sphere in both cases. However, observe that the associated filtrations are different, and such difference is taken into account by the associated eulerian magnitude homology groups.
@@
The regular path homology of the complete graph $K_{n}$ is concentrated in degree $n-1$ and is of rank the number of derangements $D(n)$.
@@
The graph $K_{n}$ is regularly diagonal. The spectral sequence converges to the homology of the flag complex of the reachability preorder, which is homotopic to the complex of injective words. By Theorem 4.2 this is, up to homotopy, a wedge of spheres in degree $n-1$ on the number of derangements of $n$. ∎
It was shown in [GLMY14] that path homology is homotopy invariant. This was generalized in [Asa23, Proposition 8.1] to the pages of the magnitude-path spectral, for $r\geq 2$. We now use Corollary 5.17 to show that regular path homology is not homotopy invariant (as opposed to path homology).
@@
Recall that complete graphs are regularly diagonal, hence the second page of the regular magnitude-path spectral sequence is non-trivial only on the diagonal $k=\ell$. By [GLMY14, Example 3.11] (see also [Asa23, Example 5.6]), the complete graph $K_{n}$ is contractible and homotopy equivalent to the one-vertex graph. If the pages of the regular magnitude-path spectral sequence were homotopy invariant, then, by restriction, also regular path homology would be homotopy invariant. However, the regular path homology of the complete graph $K_{n}$ is concentrated in degree $n-1$ and is of rank the number of derangements $D(n)$, whereas the regular path homology of the one-point graph is trivial. ∎
We conclude with some considerations on the decategorification of eulerian magnitude homology. Recall that the magnitude #$\#G(q)$ of a graph $G$ is equivalently defined as follows – see [Lei19, Definition 2.1 & Proposition 3.9]:
(8) #$\#G(q)=\sum_{k=0}^{\infty}(-1)^{k}\sum_{x_{0}\neq\dots\neq x_{k}}q^{d_{G}(x_{0},x_{1})+\dots+d_{G}(x_{k-1},x_{k})}$
@@
(9) #$\#_{r}G(q)=\sum_{k=0}^{|V(G)|}(-1)^{k}\sum_{x_{0}\neq\dots\neq x_{k}}q^{d_{G}(x_{0},x_{1})+\dots+d_{G}(x_{k-1},x_{k})}$
where the vertices $x_{0},\dots,x_{k}$ appearing in the sum are all distinct. The regular magnitude #$\#_{r}G(q)$ is not the magnitude of a graph, unless $G$ is a point. In fact, the first is a polynomial whereas the magnitude of a graph is a power series. We give a different proof which uses the (regular) magnitude-path spectral sequences, and which might shed some light on the behavior of #$\#_{r}G(q)$.
@@
By [HW17, Theorem 2.8], the magnitude of a graph $G$ agrees with the decategorification of magnitude homology of $G$. The evaluation at $q=-1$ gives the sum
@@
On the other hand, the decategorification of the eulerian magnitude homology evaluated at $q=-1$ yields the quantity
@@
Question 5.20.
Which properties does #$\#_{r}G(q)$ satisfy?

### v2

@@ 5.2. The regular magnitude-path spectral sequence
Let $T$ be a finite directed tree, a bipartite graph with alternating orientation or a transitive tournament. Then, its regular path homology is trivial.
@@
We have seen that finite directed tree, bipartite graphs with alternating orientation or transitive tournaments are regularly diagonal. Furthermore, the directed flag complexes associated to the respective reachability preorders are contractible. Hence, by Proposition 5.17, their regular path homology groups are trivial. ∎
Consider the directed graphs of Example 4.9. They are both regularly diagonal and connected, with the same number of vertices. Their regular path homology is the homology of the $2$-sphere, as $\mathrm{dFl}\circ\mathrm{Pre}$ yields the $2$-sphere in both cases. However, observe that the associated filtrations are different, and such difference is taken into account by the associated eulerian magnitude homology groups.
@@
The regular path homology of the complete graph $K_{n}$ is concentrated in degree n1n-1 and is of rank the number of derangements $D(n)$.
@@
The graph $K_{n}$ is regularly diagonal. The regular magnitude-path spectral sequence converges to the homology of the complex of injective words. By Theorem 4.2 this is, up to homotopy, a wedge of spheres in degree n1n-1 on the number of derangements of $n$. ∎
It was shown in [GLMY14] that path homology is homotopy invariant. This was generalized in [Asa23, Proposition 8.1] to the pages of the magnitude-path spectral sequence, for $r\geq 2$. We now use Corollary 5.19 to show that regular path homology is not homotopy invariant (as opposed to path homology).
@@
Recall that complete graphs are regularly diagonal, hence the second page of the regular magnitude-path spectral sequence is non-trivial only on the diagonal $k=\ell$. By [GLMY14, Example 3.11] (see also [Asa23, Example 5.6]), the complete graph $K_{n}$ is contractible and homotopy equivalent to the one-vertex graph. If the pages of the regular magnitude-path spectral sequence were homotopy invariant, then, by restriction, also regular path homology would be homotopy invariant. However, the regular path homology of the complete graph $K_{n}$ is concentrated in degree n1n-1 and is of rank the number of derangements $D(n)$, whereas the regular path homology of the one-point graph is trivial. ∎
Remark 5.21.
Chaplin, Harrington and Tillmann study in [CHT24] the homology of the directed flag complex and, by comparison with path homology, they establish a homotopy-like equivalence relation on digraph maps, under which equivalent maps induce the same morphisms on the homology of the directed flag complex. Though it may be possible to exploit their work to shed light on potential notions of homotopy invariance for the regular magnitude-path spectral sequence, pursuing this line of inquiry would require a substantial detour from the main objectives of the present paper, therefore we leave this question open.
5.3. Decategorification
We conclude with some considerations on the decategorification of eulerian magnitude homology.
Recall that the magnitude #G(q)\#G(q) of a graph $G$ is equivalently defined as follows – see [Lei19, Definition 2.1 & Proposition 3.9]:
(8) $G(q)=\sum_{k=0}^{\infty}(-1)^{k}\sum_{x_{0}\neq\dots\neq x_{k}}q^{d_{G}(x_{0},x_{1})+\dots+d_{G}(x_{k-1},x_{k})}$
@@ 5.3. Decategorification
(9) $_{r}G(q)=\sum_{k=0}^{|V(G)|}(-1)^{k}\sum_{x_{0}\neq\dots\neq x_{k}}q^{d_{G}(x_{0},x_{1})+\dots+d_{G}(x_{k-1},x_{k})}$
where the vertices $x_{0},\dots,x_{k}$ appearing in the sum are all distinct. The regular magnitude $_{r}G(q)$ is not the magnitude of a graph, unless $G$ is a point. In fact, the first is a polynomial whereas the magnitude of a graph is a power series. We give a different proof which uses the (regular) magnitude-path spectral sequences, and which might shed some light on the behavior of $_{r}G(q)$.
@@
By [HW17, Theorem 2.8], the magnitude of a graph $G$ agrees with the decategorification of magnitude homology of $G$. The evaluation at q=1q=-1 gives the sum
@@
On the other hand, the decategorification of the eulerian magnitude homology evaluated at q=1q=-1 yields the quantity
@@
Question 5.23.
Which properties does $_{r}G(q)$ satisfy?